OpenStudy (syaza97):

At what point on the curve y=tan x-x, x element (0,π/2) is the tangent line parallel to the line 2x-3y=5.

zepdrix (zepdrix):

Parallel means they have the same slope value. Do you know how to find the slope of a line which is in standard form? \(\rm 2x-3y=5\qquad\to\qquad m=?\)

OpenStudy (syaza97):

ok...i will try calculate it now..

OpenStudy (syaza97):

i got m=2/3

zepdrix (zepdrix):

So there is a line tangent to our curve y, somewhere, where the slope is parallel to this slope value.\[\large\rm y'=\frac23\]Find your derivative, then set it equal to 2/3,\[\large\rm \left(\tan x -x\right)'=\frac23\]and solve for x.

zepdrix (zepdrix):

Sorry I have to go for now.

OpenStudy (syaza97):

ok..thanks..

OpenStudy (starstudentgroup765):

GO TO SLEEP ITS 3:11 AM PS im about to also